OpenStudy (anonymous):

How would you simplify -4/4root3

OpenStudy (campbell_st):

is your question \[\frac{-4}{\sqrt[4]{3}}\]

OpenStudy (anonymous):



OpenStudy (campbell_st):

ok... so its \[\frac{-4}{4\sqrt{3}}\]

OpenStudy (campbell_st):

divide the numerator and denominator by 4

OpenStudy (anonymous):

why just 4?

OpenStudy (anonymous):

dont u have to elimate tht eroot on the bottom

OpenStudy (campbell_st):

start with that... what do you get...

OpenStudy (anonymous):

-16/root3

OpenStudy (campbell_st):

 do you know how to rationalize a denominator..?